I can't figure this out. Can anyone help? Hockey teams recieve 2 points when they win and 1 point when they tie. One season a team won a championship with 58 points. They won 11 more games than they tied. How many wins and how many ties did the team have? Thank-you for helping.
Answer by Mathtut(3670)

let W and T be the number of wins and ties respectively
:
2W+1T=58...........eq 1
W=T+11..............eq 2
:
take the value of W in eq 2 and plug it into eq 1
:
2(T+11)+T=58
:
2T+22+T=58
:
3T=36
:
T=12
:
W=T+11=12+11=23